# JSJ decompositions of knot exteriors, Dehn surgery and the $L$-space conjecture

### v3

@@ 5. Left-orderable $p/q$-surgeries on knots with $p$ small
Let $K$ be a non-trivial knot, $p=1$ or $2$, and $q\neq 0$. Then the set rational numbers $p/q$ such $K(p/q)$ is not $LO$ is contained in
@@
    $\{\varepsilon,2\varepsilon\}$ if $K=T(2,3\varepsilon)$ for some $\varepsilon=\pm 1;$ or
@@
    $\{2\varepsilon\}$ if $K$ is a $(2,\varepsilon)$-cable for some $\varepsilon=\pm 1;$ or
@@
    either {±$\{\pm 1,\pm 2\}$ or $\{\varepsilon/2,2\varepsilon/3,\varepsilon,2\varepsilon\}$ for some $\varepsilon\in\{\pm 1\}$ if $K$ is hyperbolic.
@@
In cases (1) and (2), $K(p/q)$ is not $LO$ for $p/q$ in the listed sets. On the other hand, since a hyperbolic $L$-space knot has genus at least $3$ by [24, 14], the $L$-space Conjecture predicts that for any hyperbolic knot $K$, $K(r)$ is $LO$ either for all $r\in(-\infty,5)$ or for all $r\in(-5,\infty)$. Thus it is expected that the potential exceptional slopes listed in case (3) do not arise.
@@ 5.1. $LO$ $p/q$-surgeries on satellite knots for $p$ small
Assume that $|q|\geq 2$ and that $K(p/q)$ is not $LO$, and hence is atoroidal. Let $T$ be an essential torus in the interior of $X(K)$ which bounds the exterior of a simple knot $K_{1}$ (i.e. a torus or hyperbolic knot) contained in $X(K)$. Since $K(p/q)$ is not $LO$, $T$ must compress in $K(p/q)$, and therefore as $|q|\geq 2$, [12, Theorem 2.0.1] implies that $K$ is a cable of $K_{1}$. Thus $K_{1}=K_{0}$.
Now $T$ compresses in $K(p/q)$ if and only if $|qmn-p|\leq 1$ by Theorem 2.9(2)(a). We cannot have $p=qmn$, as otherwise $|q|=1$, so $qmn=p\pm 1$. Then
@@ 5.2. $LO$ $p/q$-surgery on hyperbolic knots when $p\leq 2$
Although the question of which surgeries on a hyperbolic knot $K$ are $LO$ is wide open in general, we can deal with surgeries with $p$ small using the existence of a pseudo-Anosov flow $\Phi_{0}$ on $S^{3}\setminus K$. See [44] and [9] for background information on pseudo-Anosov flows. In the case that $p=1$ or $2$ and $K$ is hyperbolic, it is known that $K(p/q)$ is never an $L$-space. Thus we expect all such surgeries to yield manifolds that are $LO$.
Gabai and Mosher have independently shown that pseudo-Anosov flows exist on the complement of any hyperbolic link in a closed, orientable 3-manifold. More precisely, they show that given a finite depth taut foliation $\mathcal{F}$ on a compact, connected, orientable, hyperbolic $3$-manifold $M$ with non-empty boundary consisting of tori, there is a pseudo-Anosov flow on the interior of $M$ which is almost transverse to $\mathcal{F}$ (cf. [44, Theorem C(3)]). Unfortunately, no proof has been published, though Landry and Tsang have recently produced the first of several planned articles which will provide a demonstration. See [39].
@@
If $K$ is a hyperbolic knot there is a pseudo-Anosov flow $\Phi_{0}$ on its complement with degeneracy locus $\delta(\Phi_{0})$ of the form $b\mu$ or $b\mu+\lambda$ for some integer $b\neq 0$.
@@
Fix a co-oriented taut foliation $\mathcal{F}$ on $X(K_{0})$ contains a leaf that is a minimal genus Seifert surface [18]. As mentioned above, there is a pseudo-Anosov flow $\Phi_{0}$ on $S^{3}\setminus K$ which is almost transverse to $\mathcal{F}$. Then the degeneracy locus $\delta(\Phi_{0})$ of $\Phi_{0}$ is not longitudinal and it is known that $\Phi_{0}$ will extend to a pseudo-Anosov flow $\Phi_{0}(p/q)$ on $K(p/q)$ as long as the absolute value of the algebraic intersection number of $\delta(\Phi_{0})$ with the $p/q$ slope of $K$ is at least $2$ ([17]). Since $K(1/0)\cong S^{3}$ does not admit such a flow, $\delta(\Phi_{0})$ must be of the form $b\mu$ or $b\mu+\lambda$ for some non-zero integer $b$. ∎
@@
For $p=1$ or $2$, $H_{1}(K(p/q))$ is a $\mathbb{Z}/2$ vector space and if $|\delta(\Phi_{0})\cdot(p\mu+q\lambda)|\geq 2$ then $K(p/q)$ admits a pseudo-Anosov flow. The reader will verify that this inequality holds unless $bq$ is as stated in (1), (2) and (3). Corollary A.6 then shows that $\pi_{1}(K(p/q))$ is left-orderable, which completes the proof. ∎
@@
Case (1) follows from known results on left-orderable Dehn surgeries on torus knots as we stated in the beginning of §5.
For Case (3), suppose that $K$ is hyperbolic and that $K(p/q)$ is not $LO$, where $p=1$ or $2$. Theorem 5.6 implies that there is a pseudo-Anosov flow $\Phi_{0}$ on the complement of $K$ with degeneracy locus $\delta(\Phi_{0})$ of the form $b\mu$ or $b\mu+\lambda$ for some integer $b\neq 0$. If $b\geq 1$, Theorem 5.7 implies that if $p=1$ then $1/q$ is contained in either $\{-1,1\}$ or $\{1/2,1\}$. If $b\leq-1$, replace $K$ by its mirror image $K^{*}$ and note that $\Phi_{0}$ is transformed into a pseudo-Anosov flow $\Phi_{0}^{*}$ on $S^{3}\setminus K^{*}$ with degeneracy locus $\delta(\Phi_{0}^{*})$ of the form $b^{*}\mu$ or $b^{*}\mu+\lambda$ with $b^{*}=-b\geq 1$. Then as $K^{*}(-1/q)\cong K(1/q)$, $-1/q$ is contained in either $\{-1,1\}$ or $\{1/2,1\}$ and therefore $1/q$ is contained in either $\{-1,1\}$ or $\{-1,-1/2\}$.
@@
Suppose $K$ is a satellite knot $P(K_{0})$. Then $K(p/q)=P(p/q)\cup_{T}X_{0}$ where $X_{0}$ is the exterior of $K_{0}$ and $T=\partial P(p/q)=\partial X_{0}$. Since $|H_{1}(K(p/q))|=p$, if $T$ is incompressible in $P(p/q)$ then $K(p/q)$ is $LO$ by Theorem 5.3.
So assume that $T$ compresses in $P(p/q)$. Then by Theorem 2.9 either (i) $P=C_{m,n}$ and $p/q$ is the cabling slope $mn$, or (ii) $P(p/q)$ is a solid torus.
Possibility (i) gives case (2) of the theorem. So assume (ii) holds. Then either (a) $P=C_{m,n}$ and $\Delta(p/q,mn/1)=1$, or (b) $P$ is a 1-bridge braid and $|q|=1$. It is easy to verify that the only solutions to (a) are $p=1,q=\varepsilon,m=2,n=\varepsilon$, and $p=2,q=\varepsilon,m=3,n=\varepsilon$, for some $\varepsilon=\pm 1$. Thus in both cases (a) and (b), $|q|=1$. Moreover, in case (a) $w=w(P)=m$, and in case (b) $w\geq 4$.
We have $K(p/q)=K(p)=K_{0}(p/w^{2})$. If $K_{0}$ is a torus knot then $K_{0}(p/w^{2})$ is $LO$ since $p/w^{2}<1$. The same conclusion holds if $K_{0}$ is hyperbolic by Case (3). Finally, the same conclusion holds if $K_{0}$ is a satellite knot by the argument above that any non-integral surgery $p/q$ on a satellite knot with $p=1$ or $2$ is $LO$. ∎
@@
Let $K$ be a hyperbolic knot such that some $n$-fold cyclic branched cover of $K$, $n\geq 2$, is not $LO$. Then $K(1/q)$ and $K(2/q)$ are $LO$ for $|q|\geq 2$. ∎
@@
Note that Corollaries 5.8 and 5.10 can also be proven using the universal circle actions (see [11, 5]) associated to the taut foliations on the surgered manifolds constructed by Roberts in [50] and [49] respectively. The Euler class of these universal circle actions vanishes by [34, Proposition 2.1] and therefore the fundamental groups of the surgered manifolds are LO (cf. [5, Theorem 8.3]).
For general $p/q\in\mathbb{Q}$, we can find a pseudo-Anosov flow $\Phi(p/q)$ on $K(p/q)$ as long as $|\delta(\Phi)\cdot(p\mu+q\lambda)|\geq 2$ and consequently faithful representations $\rho_{p/q}:\pi_{1}(K(p/q))\to\mbox{Homeo}_{+}(S^{1})$ (cf. the Appendix). As noted in the Appendix, $\pi_{1}(K(p/q))$ is left-orderable if the Euler class $e(\rho_{p/q})\in H^{2}(K(p/q))$ vanishes. Equivalently, the Euler class of the normal bundle to the flow $e(\nu_{\Phi(p/q)})$ vanishes (Proposition A.4). But this does not appear to happen very frequently. For instance, though the methods of [34] can be often be used to show that there are infinitely many values of $p/q$ for which $e(\rho_{p/q})=0$, that paper shows that the set of such $p/q$ is typically nowhere dense in the reals.

### v4

@@ 5. Left-orderable $p/q$-surgeries on knots with $p$ small
Let $K$ be a non-trivial knot and let $E$ denote the set of rational numbers $p/q$ such that $p=1$ or $2$, $q\neq 0$, and $K(p/q)$ is not $LO$. If $E$ is non-empty, then one of the following occurs:
@@
    $K=T(2,3\varepsilon)$ for some $\varepsilon=\pm 1$ and $E=\{\varepsilon,2\varepsilon\};$ or
@@
    $K$ is a $(2,\varepsilon)$-cable for some $\varepsilon=\pm 1$ and $E=\{2\varepsilon\};$ or
@@
    $K$ is hyperbolic and there is some $\varepsilon\in\{\pm 1\}$ such that $E$ is contained in either $\{\varepsilon,2\varepsilon\}$ or $\{\varepsilon/2,2\varepsilon/3,\varepsilon,2\varepsilon\}$.
@@
Since a hyperbolic $L$-space knot has genus at least $3$ by [24, 14], the $L$-space Conjecture predicts that for any hyperbolic knot $K$, $K(r)$ is $LO$ either for all $r\in(-\infty,5)$ or for all $r\in(-5,\infty)$. Thus it is expected that the potential exceptional slopes listed in case (3) of the theorem do not arise.
@@ 5.1. $LO$ $p/q$-surgeries on satellite knots for $p$ small
Assume that $|q|\geq 2$ and that $K(p/q)$ is not $LO$, and hence is atoroidal. Then $T$, the boundary of the exterior of $K_{0}$, must compress in $C_{m,n}(p/q)$. By Theorem 2.9(2)(a), and since $|q|\geq 2$, this happens if and only if $qmn=p\pm 1$. Then
@@ 5.2. $LO$ $p/q$-surgery on hyperbolic knots when $p\leq 2$
The question of which surgeries on a hyperbolic knot $K$ are $LO$ is wide open in general. In the case that $p=1$ or $2$, $q\neq 0$, and $K$ is hyperbolic, it is known that $K(p/q)$ is never an $L$-space. Thus we expect all such surgeries to yield manifolds that are $LO$. In this section, we will present some results regarding the left-orderability of the fundamental groups of these manifolds and describe their proofs using known results on pseudo-Anosov flows. See [44] and [11, §6.6] for background information on pseudo-Anosov flows.
A pseudo-Anosov flow $\Phi_{0}$ on the complement of a knot $K$ determines an essential lamination in the interior of $M$, i.e., the (un)-stable lamination of $\Phi_{0}$. The connected complementary component $C$ of the lamination that contains $\partial M$ is a bundle over $S^{1}$ whose fiber is an ideal polygon with an open disk removed from its interior. It is easy to see that this component is homeomorphic to $\partial M\times[0,1]$ with a collection of simple closed curves removed on $\partial M\times\{1\}$, where $\partial M$ is identified with $\partial M\times\{0\}$. This collection of simple closed curves is called the degeneracy locus of the pseudo-Anosov flow. See [11, Definition 6.42].
Fried described an operation which extends a pseudo-Anosov flow $\Phi_{0}$ on the complement of a hyperbolic knot $K$ to the closed manifold obtained by $p/q$-surgery along $K$ [17]. The condition for this to apply is given in terms of the degeneracy locus:
Theorem 5.6 (Goodman, Fried).
A pseudo-Anosov flow $\Phi_{0}$ on the complement of $K$ will extend to a pseudo-Anosov flow $\Phi_{0}(p/q)$ on $K(p/q)$ as long as the absolute value of the algebraic intersection number of $\delta(\Phi_{0})$ with the $p/q$ slope of $K$ is at least $2$.
Gabai and Mosher have independently shown that pseudo-Anosov flows exist on the complement of any hyperbolic link in a closed, orientable 3-manifold. More precisely, they show that given a finite depth taut foliation $\mathcal{F}$ on a compact, connected, orientable, hyperbolic $3$-manifold $M$ with non-empty boundary consisting of tori, there is a pseudo-Anosov flow on the interior of $M$ which is almost transverse to $\mathcal{F}$ (cf. [44, Theorem C(3)]). Unfortunately, no proof has been published, though Landry and Tsang have recently produced the first of several planned articles which will provide a demonstration. See [39]. The following fact about the degeneracy loci of these flows is well known to experts. We include a proof for completeness.
@@
Let $K$ be a hyperbolic knot in the $3$-sphere. There is a pseudo-Anosov flow $\Phi_{0}$ on the complement of $K$ with degeneracy locus $\delta(\Phi_{0})$ of the form $b\mu$ or $b\mu+\lambda$ for some integer $b\neq 0$.
@@
By [18], there is a finite depth foliation $\mathcal{F}$ on $X(K)$ whose depth zero leaf is a minimal genus Seifert surface for $K$. Then by the results of Gabai and Mosher discussed above, there exists a pseudo-Anosov flow on the interior of $X(K)$, denoted by $\Phi_{0}$, that is almost transverse to $\mathcal{F}$. Since $K(1/0)\cong S^{3}$ has trivial fundamental group, it supports no pseudo-Anosov flow. Theorem 5.6 then implies that $\delta(\Phi_{0})$ must be of the form $b\mu$ or $b\mu+\lambda$ for some integer $b$. In the latter case, $b\neq 0$ since the flow is almost transverse to the foliation, in particular the depth zero leaf  [^6] See [11, Definition 6.45] for the definition of a pseudo-Anosov flow being almost transverse to a foliation. . ∎
@@
For $p=1$ or $2$, $H_{1}(K(p/q))$ is a $\mathbb{Z}/2$ vector space and if $|\delta(\Phi_{0})\cdot(p\mu+q\lambda)|\geq 2$ then Theorem 5.6 implies that $K(p/q)$ admits a pseudo-Anosov flow. The reader will verify that this inequality holds unless $bq$ is as stated in (1), (2) and (3). Corollary A.6 then shows that $\pi_{1}(K(p/q))$ is left-orderable, which completes the proof. ∎
@@
Suppose that $K$ is a non-trivial knot in $S^{3}$, $p=1$ or $2$, and $q\neq 0$, and $K(p/q)$ is not $LO$. As a non-trivial knot, $K$ is either a torus knot, a hyperbolic knot, or a satellite knot. We consider these three cases separately.
The known results on left-orderable Dehn surgeries on torus knots stated at the beginning of §5 imply that $K=T(2,3\varepsilon)$ for some $\varepsilon\in\{\pm 1\}$ and $p/q\in\{\varepsilon,2\varepsilon\}$. This yields case (1) of the theorem.
Suppose next that $K$ is hyperbolic. Theorem 5.6(2) implies that there is a pseudo-Anosov flow $\Phi_{0}$ on the complement of $K$ with degeneracy locus $\delta(\Phi_{0})$ of the form $b\mu$ or $b\mu+\lambda$ for some integer $b\neq 0$. If $b\geq 1$, Theorem 5.8 implies that if $p=1$ then $1/q$ is contained in either $\{-1,1\}$ or $\{1/2,1\}$. If $b\leq-1$, replace $K$ by its mirror image $K^{*}$ and note that $\Phi_{0}$ is transformed into a pseudo-Anosov flow $\Phi_{0}^{*}$ on $S^{3}\setminus K^{*}$ with degeneracy locus $\delta(\Phi_{0}^{*})$ of the form $b^{*}\mu$ or $b^{*}\mu+\lambda$ with $b^{*}=-b\geq 1$. Then as $K^{*}(-1/q)\cong K(1/q)$, $-1/q$ is contained in either $\{-1,1\}$ or $\{1/2,1\}$ and therefore $1/q$ is contained in either $\{-1,1\}$ or $\{-1,-1/2\}$.
@@
Finally suppose that $K$ is a satellite knot $P(K_{0})$ and that $K(p/q)$ is not $LO$. Then $K(p/q)=P(p/q)\cup_{T}X_{0}$ where $X_{0}$ is the exterior of $K_{0}$ and $T=\partial P(p/q)=\partial X_{0}$. Since $|H_{1}(K(p/q))|=p$, if $T$ is incompressible in $P(p/q)$ then $K(p/q)$ would be $LO$ by Theorem 5.3, contrary to our hypotheses.
Thus $T$ compresses in $P(p/q)$. Then by Theorem 2.9 either (i) $P=C_{m,n}$ and $p/q$ is the cabling slope $mn$, or (ii) $P(p/q)$ is a solid torus.
If possibility (i) arises, then $q=\varepsilon\in\{\pm 1\}$ and therefore $mn=\varepsilon p$. Since $m\geq 2$, we must have $p=m=2$ and $n=\varepsilon$. Conversely, when $p,q,m,n$ take on these values, $K(p/q)$ has an $\mathbb{R}P^{3}$ summand, so is not $LO$. To complete the proof of Case (2) we must rule out possibility (ii).
Assume that (ii) holds. Then Theorems 2.9 and 2.13 imply that either (a) $P=C_{m,n}$ and $\Delta(p/q,mn/1)=1$, or (b) $P$ is a 1-bridge braid and $|q|=1$. It is easy to verify that the only solutions to (a) are $p=1,q=\varepsilon,m=2,n=\varepsilon$, and $p=2,q=\varepsilon,m=3,n=\varepsilon$, for some $\varepsilon=\pm 1$. Thus in both cases (a) and (b), $|q|=1$. Moreover, in case (a) $w=w(P)=m\geq 2$, and in case (b) $w\geq 4$ by Lemma 2.10.
We have $K(p/q)=K(p)=K_{0}(p/w^{2})$ by Theorem 2.13, where $w^{2}\geq 4$. If $K_{0}$ is a torus knot then $K_{0}(p/w^{2})$ is $LO$ since $p/w^{2}<1$. The same conclusion holds if $K_{0}$ is hyperbolic by Case (3). Finally, the conclusion holds if $K_{0}$ is a satellite knot by the argument above that any non-integral surgery $p/q$ on a satellite knot with $p=1$ or $2$ is $LO$. ∎
@@
Let $K$ be a hyperbolic knot such that some $n$-fold cyclic branched cover of $K$, $n\geq 2$, is not $LO$. Then $K(1/q)$ and $K(2/q)$ are $LO$ for $|q|\geq 2$.
@@
Note that Corollaries 5.9 and 5.11 can also be proven using the universal circle actions (see [12, 5]) associated to the taut foliations on the surgered manifolds constructed by Roberts in [50] and [49] respectively. The Euler class of these universal circle actions vanishes by [34, Proposition 2.1] and therefore the fundamental groups of the surgered manifolds are LO (cf. [5, Theorem 8.3]).
For general $p/q\in\mathbb{Q}$, we can find a pseudo-Anosov flow $\Phi(p/q)$ on $K(p/q)$ as long as $|\delta(\Phi)\cdot(p\mu+q\lambda)|\geq 2$ and consequently faithful representations $\rho_{p/q}:\pi_{1}(K(p/q))\to\mbox{Homeo}_{+}(S^{1})$ (see the Appendix). As noted in the Appendix, $\pi_{1}(K(p/q))$ is left-orderable if the Euler class $e(\rho_{p/q})\in H^{2}(K(p/q))$ vanishes. Equivalently, the Euler class of the normal bundle to the flow $e(\nu_{\Phi(p/q)})$ vanishes (Proposition A.4). But this does not appear to happen very frequently. For instance, though the methods of [34] can be often be used to show that there are infinitely many values of $p/q$ for which $e(\rho_{p/q})=0$, they also show that the set of such $p/q$ is typically nowhere dense in the reals.